OpenStudy (anonymous):

someone help me with a few geometry questions please <3

OpenStudy (anonymous):

i need to solve 3x + 5 = 10x +2 5 = 7x + 2 idk what to do from here someone please help

OpenStudy (waynex):

Subtract 2 from both sides, then divide both sides by seven. Try that. If that doesn't work, let me know.

OpenStudy (anonymous):

ill post the question bc im so lost im trying to catch up on my back work i have till midnight east time and im just getting irritated.

OpenStudy (anonymous):

5. In the square ABCD, AE = 3x + 5 and BD = 10x + 2. What is the length of AC ? 2 11 22 32

OpenStudy (anonymous):

i get stuck on these stupid problems all the time.

jimthompson5910 (jim_thompson5910):

the diagonals are congruent, so AC = BD furthermore, each diagonal cuts the other diagonal in half, this means AE = (1/2)*AC which rearranges to 2*AE = AC since AC = BD, we can say 2*AE = BD

jimthompson5910 (jim_thompson5910):

now plug in the given expressions AE = 3x+5, BD = 10x+2 2*AE = BD 2*(3x+5) = 10x+2 6x + 10 = 10x + 2 ... distribute 10 = 4x + 2 ... Subtract 6x from both sides. 8 = 4x ... Subtract 2 from both sides. I'll let you finish

OpenStudy (anonymous):

okay so x = 2, plugging that in for BD which = AC 10(2)+2= 4... but thats not an option ugh

jimthompson5910 (jim_thompson5910):

how does 10*2+2 equal 4?

jimthompson5910 (jim_thompson5910):

you just made a silly math error, but that's ok

OpenStudy (anonymous):

wait no, 3(2)+5 = 11 nvm lol and i meant 40 but i got it now thank you. can you help me with a couple more?

jimthompson5910 (jim_thompson5910):

hopefully you got AC = BD AC = 10x+2 AC = 10*2+2 AC = 20+2 AC = 22

jimthompson5910 (jim_thompson5910):

yeah i can help with a few more

OpenStudy (anonymous):

okay thank you this ones similar, 5x-5=3x+3 5x=3x+7 ... i keep getting decimal numbers, blah

jimthompson5910 (jim_thompson5910):

5x-5=3x+3 5x = 3x + 8 ... Add 5 to both sides. 2x = 8 ... Subtract 3x from both sides. x = ???

jimthompson5910 (jim_thompson5910):

you just made the error in thinking that 5+3 was 7 when it should have been 8

OpenStudy (anonymous):

oh okay well i didnt know since i took the -5 and brought it to the other side if i could take the 3x to the oppisite side. so x=4 right?

jimthompson5910 (jim_thompson5910):

you got it, the solution is x = 4

jimthompson5910 (jim_thompson5910):

3x is a number like -5 is, so you can add/subtract them to both sides if necessary

OpenStudy (anonymous):

okay i will have a few here eventually im trying to do most of them myself, will you be on here for a little while yet?

jimthompson5910 (jim_thompson5910):

yes I'll be on if you need more help, but see how far you get

OpenStudy (anonymous):

the answer to #5 is actually 22, not 11. i just got that answer correct.